# Online LLM watermark detection via e-processes

### v1

@@ 4.2 Fixed calibrators with adaptive weights
We first consider some fixed choices of deterministic calibrators, which are nonnegative decreasing functions on $[0,1]$ that integrate to at most $1$. Some common examples (Vovk and Wang,, 2021) are, as functions of $p$,
@@
For better power, $(\lambda_{t})_{t\in\mathbb{T}}$ should be chosen adaptively. Using the idea of the empirically adaptive e-process in Ramdas and Wang, (2025, Section 7.8), $\lambda_{t}$ is given by $\lambda_{1}=0$ and
@@ 4.3 Online Grenander e-processes
Let $\mathcal{D}$ be the set of all decreasing density functions on $[0,1]$ (that is, the set of calibrators). Define the online Grenander (OG) calibrator (Hore et al.,, 2026), by
@@
where we set $Y_{0}=0$. The computation of (14) is straightforward as it coincides with the classic Grenander, (1956) estimator applied to $(Y_{0},Y_{1},\dots,Y_{t-1})$: It is known that the unique maximizer in (14) $f^{\mathrm{OG}}_{t}$, which is unique on $(0,1]$ and flexible at 0, is the left derivative of the smallest concave majorant of the empirical distribution function of $Y_{0},Y_{1},\dots,Y_{t-1}$ (see e.g., Chapter 8 of Devroye,, 1987). We say that it is unique because the value of $f_{t}^{\mathrm{OG}}$ at 0 does not matter, and we can safely set $f_{t}^{\mathrm{OG}}$ to be continuous at 0. The function $f^{\mathrm{OG}}_{t}$ is a left-continuous step function. Let
@@
The OG calibrator is useful in several settings. For instance, Hore et al., (2026) showed that in an iid setting, the OG e-process with range $a,b]\subseteq(0,\infty)$ has an optimal regret bound.
@@ 4.4 Average e-processes
The test based on the average e-process in (17) may have better performance than both tests based on the individual ones, as shown by our empirical results in Section 6. As a feature of e-processes, the arithmetic average of two e-processes typically has the advantages of the two; see e.g., the mixture e-processes in Ramdas and Wang, (2025). In particular, it is straightforward to see that the asymptotic growth rate of the average e-process is the maximum of those of the two individual e-processes. Taking an average of more than two e-processes (e.g., with different calibrators) also has such a feature, but it does not offer visible further improvements on power performance according to our empirical experiments.
@@ 5 Theoretical power guarantees
Note that the distribution of $(\mathbf{P}_{t})_{t\in\mathbb{T}}$ is the same under both the null and the alternative hypotheses. The assumption $\delta>0$ means that the temperature of the LLM is bounded away from 0, and thus the token generating process has enough randomness. The assumption of $m$-dependence means that the generation of one token only relies a maximum of $m$ past tokens. This assumption is needed to apply a form of the law of large numbers. Some other forms of weak or short-term dependence such as $\alpha$-mixing are also sufficient for this purpose, and we omit such discussions.
@@
Since $1-\lambda+\lambda E_{t}$, $t\in\mathbb{T}$ are identically distributed under $\mathbb{P}$ with finite mean, $\mathbb{E}^{\mathbb{P}}[(\log(1-\lambda+\lambda E_{t}))^{2}]$ is bounded above across $t$. Hence, $\mathbb{E}^{\mathbb{Q}}[(\log(1-\lambda+\lambda E_{t}))^{2}]$ is also bounded above. This allows us to use the strong law of large numbers for $m$-dependent sequences (e.g., Theorem 1 of Hansen,, 1991), which gives
@@
where the inequality follows from Chebyshev’s association inequality, using that $g$ is nonconstant. This yields, in particular, $\mathbb{E}^{\mathbb{Q}^{\mathbf{P}^{*}}}[g(1-Y)]>1$. By using Theorem 3.14 of Ramdas and Wang, (2025), this further implies
@@
First, consider the weight-adaptive e-process $M$ in (13). For $t\in\mathbb{T}$, let $E_{t}=g(1-Y_{t})$. Assumption 1 (ii) guarantees $\mathbb{E}^{\mathbb{Q}}[E_{t}]>1$. Let
@@
Note that the optimizer $\lambda^{*}$ is unique because of the strict concavity of the log function, $\lambda^{*}$ does not depend on $t$ because $E_{1},E_{2},\dots$ are iid, and $\lambda^{*}>0$, which follows from Theorem 3.14 of Ramdas and Wang, (2025) and the condition $\mathbb{E}^{\mathbb{Q}}[E_{t}]>1$. Note that
@@
Proposition EC.2 of Wang et al., (2025) gives
@@
To show the statement on the OG e-process $M$ in (16) with range $[a,b]$, $0<a<1<b<\infty$, we use Theorem 3.1 of Hore et al., (2026), which gives
@@ 6 Experiments on open-source LLMs
  • The sum-based method using the score function $h_{\rm ars}(y)=-\log(1-y)$ (Aaronson,, 2023).
  • The sum-based method using the score function $h_{\log}(y)=\log(y)$ (Fernandez et al.,, 2023; Kuditipudi et al.,, 2024).
  • The optimized sum-based methods using the score function $h_{\text{gum},d}^{*}$ proposed by Li et al., 2025b , with the hyper-parameter $d=0.1$ and $d=0.01$, respectively.
@@
We first evaluate these methods on simulated data. The experimental setup essentially follows the approach described in Li et al., 2025b . We generate a vocabulary $\mathcal{W}$ of size 1,000 and assess Type II, Type I and sequential Type I errors throughout 1,000 repeated experiments. Here, sequential Type I errors are computed through sequential monitoring; that is, for $t=1,2,\dots,$ we run the test on $t$ tokens, and declare rejection (watermark is detected) as soon as the test for $t$ tokens rejects. For traditional tests that are designed for fixed token length, this will inflate the Type I error.
In our simulation, we model LLM’s NTP distributions as spike distributions: we set its largest probability as $1-\Delta_{t}$, where $\Delta_{t}$ is iid sampled from $\mathrm{U}(0.001,\delta)$ for $t=1,\cdots,700$ , and the remaining probabilities are firstly sampled from $\mathrm{U}(0,1)$ and subsequently normalized such that $\mathbf{P}_{t}||_{1}=1$. We assess (sequential) Type I error rates on unwatermarked text, and Type II error rates on text generated from the same model with the Gumbel-max watermark. The error rate curves are depicted in Figure 1.
We proceed to investigate the empirical performance of watermark detection methods for text sequences generated by the LLM OPT-1.3B, following the experimental setup in prior works (Kirchenbauer et al.,, 2023; Li et al., 2025b, ). Although OPT-1.3B has fewer parameters compared to commercial LLMs (and it may therefore be regarded as a small language model), evaluating watermark detection methods is a statistical task that is not about the quality or intelligence level of the text generated; we expect similar patterns for the power of watermark detection for more complicated LLMs. The average Type I and Type II error curves are evaluated based on 500 independent repeated experiments.
@@
    While Type II error rates for all methods tend to decrease as the text sequences grow longer, sum-based methods exhibit a slight inflation in Type II error rates under low-temperature settings. This inflation in sum-based methods likely stems from the emergence of degenerate NTPs at lower temperatures, particularly in extended sequences—a phenomenon documented in recent studies (Peeperkorn et al.,, 2024; Li et al., 2025a, ). Since tokens generated from degenerate distributions are deterministic and independent of the watermark’s pseudo-randomness, they act as noise that interferes with sum-based detection. In contrast, e-process methods maintain a consistent and monotonic reduction in Type II errors, making them suitable in such situations.
@@ 7 Conclusion
In this paper, we presented a unified framework for LLM watermark detection based on e-processes, providing anytime-valid guarantees both theoretically and empirically. Various empirically adaptive e-process construction methods are developed to improve the testing power. In addition, our results analyze the power properties of the proposed procedures, offering insights into their statistical efficiency and asymptotic behavior. Experimental evaluations further demonstrate the advantages of the proposed framework over existing watermark detection methods.
The present study remains preliminary in several aspects. The theoretical characterization of optimal e-process constructions is not yet complete, especially with respect to achieving the upper bounds of e-power. Furthermore, the selection and design of statistically efficient pivotal statistics tailored to different watermark mechanisms warrant deeper investigation. Future work may focus on refining the theoretical limits of sequential testing power within this framework, exploring adaptive and data-driven construction strategies, and developing principled criteria for selecting optimal test statistics under varying watermark schemes. These directions are expected to further strengthen the statistical foundations and practical effectiveness of LLM watermark detection.

### v2

@@ 4.2 Fixed calibrators with adaptive weights
We first consider some fixed choices of deterministic calibrators, which are nonnegative decreasing functions on $[0,1]$ that integrate to at most $1$. Some common examples (Vovk and Wang, 2021) are, as functions of $p$,
@@
For better power, $(\lambda_{t})_{t\in\mathbb{T}}$ should be chosen adaptively. Using the idea of the empirically adaptive e-process in Ramdas and Wang (2025, Section 7.8), $\lambda_{t}$ is given by $\lambda_{1}=0$ and
@@ 4.3 Online Grenander e-processes
Let $\mathcal{D}$ be the set of all decreasing density functions on $[0,1]$ (that is, the set of calibrators). Define the online Grenander (OG) calibrator (Hore et al., 2026), by
@@
where we set $Y_{0}=0$. The computation of (15) is straightforward as it coincides with the classic Grenander (1956) estimator applied to $(Y_{0},Y_{1},\dots,Y_{t-1})$: It is known that the unique maximizer in (15) $f^{\mathrm{OG}}_{t}$, which is unique on $(0,1]$ and flexible at 0, is the left derivative of the smallest concave majorant of the empirical distribution function of $Y_{0},Y_{1},\dots,Y_{t-1}$ (see e.g., Chapter 8 of Devroye, 1987). We say that it is unique because the value of $f_{t}^{\mathrm{OG}}$ at 0 does not matter, and we can safely set $f_{t}^{\mathrm{OG}}$ to be continuous at 0. The function $f^{\mathrm{OG}}_{t}$ is a left-continuous step function. Let
@@
The OG calibrator is useful in several settings. For instance, Hore et al. (2026) showed that in an iid setting, the OG e-process with range $a,b]\subseteq(0,\infty)$ has an optimal regret bound.
@@ 4.4 Average e-processes
The test based on the average e-process in (18) may have better performance than both tests based on the individual ones, as shown by our empirical results in Section 6. As a feature of e-processes, the arithmetic average of two e-processes typically has the advantages of the two; see e.g., the mixture e-processes in Ramdas and Wang (2025). In particular, it is straightforward to see that the asymptotic growth rate of the average e-process is the maximum of those of the two individual e-processes. Taking an average of more than two e-processes (e.g., with different calibrators) also has such a feature, but it does not offer visible further improvements on power performance according to our empirical experiments.
@@ 5 Theoretical power guarantees
Note that the distribution of $(\mathbf{P}_{t})_{t\in\mathbb{T}}$ is the same under both the null and the alternative hypotheses. The assumption $\delta>0$ means that the temperature of the LLM is bounded away from 0, and thus the token generating process has enough randomness. The assumption of $m$-dependence means that the generation of one token only relies on a maximum of $m$ past tokens. This assumption is needed to apply a form of the law of large numbers. Some other forms of weak or short-term dependence such as $\alpha$-mixing are also sufficient for this purpose, and we omit such discussions.
@@
Since $1-\lambda+\lambda E_{t}$, $t\in\mathbb{T}$ are identically distributed under $\mathbb{P}$ with finite mean, $\mathbb{E}^{\mathbb{P}}[(\log(1-\lambda+\lambda E_{t}))^{2}]$ is bounded above across $t$. Hence, $\mathbb{E}^{\mathbb{Q}}[(\log(1-\lambda+\lambda E_{t}))^{2}]$ is also bounded above. This allows us to use the strong law of large numbers for $m$-dependent sequences (e.g., Theorem 1 of Hansen, 1991), which gives
@@
where the inequality follows from Chebyshev’s association inequality, using that $g$ is nonconstant. This yields, in particular, $\mathbb{E}^{\mathbb{Q}^{\mathbf{P}^{*}}}[g(1-Y)]>1$. By using Theorem 3.14 of Ramdas and Wang (2025), this further implies
@@
First, consider the weight-adaptive e-process $M$ in (14). For $t\in\mathbb{T}$, let $E_{t}=g(1-Y_{t})$. Assumption 4 guarantees $\mathbb{E}^{\mathbb{Q}}[E_{t}]>1$. Let
@@
Note that the optimizer $\lambda^{*}$ is unique because of the strict concavity of the log function, $\lambda^{*}$ does not depend on $t$ because $E_{1},E_{2},\dots$ are iid, and $\lambda^{*}>0$, which follows from Theorem 3.14 of Ramdas and Wang (2025) and the condition $\mathbb{E}^{\mathbb{Q}}[E_{t}]>1$. Note that
@@
Proposition EC.2 of Wang et al. (2025) gives
@@
To show the statement on the OG e-process $M$ in (17) with range $[a,b]$, $0<a<1<b<\infty$, we use Theorem 3.1 of Hore et al. (2026), which gives
@@ 6 Experiments on open-source LLMs
  • The sum-based method using the score function $h_{\rm ars}(y)=-\log(1-y)$ (Aaronson, 2023).
  • The sum-based method using the score function $h_{\log}(y)=\log(y)$ (Fernandez et al., 2023; Kuditipudi et al., 2024).
  • The optimized sum-based methods using the score function $h_{\text{gum},d}^{*}$ proposed by Li et al. (2025b), with the hyper-parameter $d=0.1$ and $d=0.01$, respectively.
@@
We first evaluate these methods on simulated data. The experimental setup essentially follows the approach described in Li et al. (2025b). We generate a vocabulary $\mathcal{W}$ of size 1,000 and assess Type II, Type I and sequential Type I errors throughout 1,000 repeated experiments. Here, sequential Type I errors are computed through sequential monitoring; that is, for $t=1,2,\dots,$ we run the test on $t$ tokens, and declare rejection (watermark is detected) as soon as the test for $t$ tokens rejects. For traditional tests that are designed for fixed token length, this will inflate the Type I error.
In our simulation, we model LLM’s NTP distributions as spike distributions: we set its largest probability as $1-\Delta_{t}$, where $\Delta_{t}$ is iid sampled from $\mathrm{U}(0.001,\delta)$ for $t=1,\dots,700$, and the remaining probabilities are firstly sampled from $\mathrm{U}(0,1)$ and subsequently normalized such that $\mathbf{P}_{t}||_{1}=1$. We assess (sequential) Type I error rates on unwatermarked text, and Type II error rates on text generated from the same model with the Gumbel-max watermark. The error rate curves are depicted in Figure 1.
We proceed to investigate the empirical performance of watermark detection methods for text sequences generated by the LLM OPT-1.3B, following the experimental setup in prior works (Kirchenbauer et al., 2023; Li et al., 2025b). Although OPT-1.3B has fewer parameters compared to commercial LLMs (and it may therefore be regarded as a small language model), evaluating watermark detection methods is a statistical task that is not about the quality or intelligence level of the text generated; we expect similar patterns for the power of watermark detection for more complicated LLMs. The average Type I and Type II error curves are evaluated based on 500 independent repeated experiments.
@@
    While Type II error rates for all methods tend to decrease as the text sequences grow longer, sum-based methods exhibit a slight inflation in Type II error rates under low-temperature settings. This inflation in sum-based methods likely stems from the emergence of degenerate NTPs at lower temperatures, particularly in extended sequences—a phenomenon documented in recent studies (Peeperkorn et al., 2024; Li et al., 2025a). Since tokens generated from degenerate distributions are deterministic and independent of the watermark’s pseudo-randomness, they act as noise that interferes with sum-based detection. In contrast, e-process methods maintain a consistent and monotonic reduction in Type II errors, making them well suited to such situations.
@@ 7 Conclusion
In this paper, we presented a unified framework for LLM watermark detection based on e-processes, providing anytime-valid guarantees both theoretically and empirically. We show in Theorem 2 that our proposed method is the only admissible and unbiased method for this testing problem under some mild assumptions, satisfied by, e.g., the Gumbel-max watermark. Various empirically adaptive e-process construction methods are developed to improve the testing power. In addition, our results analyze the power properties of the proposed procedures, illustrating their statistical efficiency and asymptotic behavior. Experimental evaluations further demonstrate the advantages of the proposed framework over existing watermark detection methods.
The present study remains preliminary in several aspects. The theoretical characterization of optimal e-process constructions is not yet complete, especially with respect to achieving the theoretical upper bounds on the e-power. Furthermore, the selection and design of statistically efficient pivotal statistics tailored to different watermark mechanisms warrant deeper investigation. Future work may focus on refining the theoretical limits of sequential testing power within this framework, exploring adaptive and data-driven construction strategies, and developing principled criteria for selecting optimal test statistics under varying watermark schemes. These directions are expected to further strengthen the statistical foundations and practical effectiveness of LLM watermark detection.